OpenStudy (anonymous):

The perimeter of a rectangle is 48 cm. The length is 6 cm less than twice the width. Which system would model this situation, and what are the length and the width?

OpenStudy (cwrw238):

let w = width, l = length then: 2w + 2l = 48 and l = 2w - 6 so we have the system 2w + 2l = 48 2w - l = 6